OpenStudy (anonymous):

Find the average of the function over the given interval and all values of x in the interval for which the function equals its average value. (Round your answer to three decimal places.) f(x)=4 cos(x), [0,pi/6] (x,y)=(___?___)

OpenStudy (dumbcow):

the avg value is: \[\frac{1}{b-a} \int\limits_a^b f(x)\] \[\rightarrow \frac{6}{\pi}\int\limits _0 ^{\pi/6} 4\cos x dx\]

OpenStudy (anonymous):

Not sure where to go from there

OpenStudy (anonymous):

compute the integral, which is not too hard because the anti derivative of \(\cos(x)\) is \(\sin(x)\)

OpenStudy (anonymous):

you have \[\frac{24}{\pi}\sin(\frac{\pi}{6})\] \[=\frac{24}{\pi}\times \frac{1}{2}=\frac{12}{\pi}\]

OpenStudy (anonymous):

not sure how you are going to solv e \[4\cos(x)=\frac{12}{\pi}\] i guess divide by 4 and get \[\cos(x)=\frac{3}{\pi}\] so \[x=\cos^{-1}\left(\frac{3}{\pi}\right)\]